[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Apr 20, 2017

Craig Mattocks - NOAA Federal craig.mattocks at noaa.gov
Thu Apr 20 15:51:39 UTC 2017


Additional issues:

1. Intermittent network problems at the NOAA Web Operation Center (WOC) in Boulder, CO are causing many NOAA web sites, including NHC’s, to fail to load at times. This might have been triggered by the failed-and-rolled-back College Park Data Center switch upgrade (RFC #3039) on Monday.

2. A power outage occurred yesterday afternoon at the GOES-16 Consolidated Backup (CBU) facility in Fairmont, West Virginia. This may partially explain the loss of GOES-16 imagery in N-AWIPS. Power has since been restored.

HSU:

1. ATCF issues: Hurricane Specialists had to once again submit HWRF/GFDL CARQ twice at 06Z in order for the SDM to receive it. These seems to be an intermittent network connection problem between atcfsvr1 and WCOSS. TSB will monitor.

2. N-AWIPS issues: When calling up Data, the Accept and Close buttons do not appear when loading the Misc data type. TSB will investigate.

Craig


> On Apr 20, 2017, at 10:30 AM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:
> 
> Here is today's summary of NHC computer operations: 
> --- Craig
> 
> ----------
> 
> TAFB:
> 1. Security camera display in MFL and TAFB now missing two cameras now. An NCO Help Desk ticket will be submitted.
> 
> HSU:
> 
> Other items:
> 1. The GFS upgrade has been approved by the NCEP Director. Implementation is now planned for July 2017.
> 
> 2. The GFDL Hurricane model will be discontinued effective with the GFS upgrade in July 2017. See the Service Change Notice (SCN) at:
> http://www.nws.noaa.gov/os/notification/scn17-45discontinue_ghm_v2.htm
> 
> 3. NESDIS is doing navigation tests on the GOES-16 mesoscale windows. Memory leaks have been discovered in the ISatSS processing code which is causing the latency and data gaps in the GOES-16 feed noted by HSU and TAFB forecasters.
> 
> 4. GOES-16 planned maintenance:
> a) Tests of mesoscale domains not being available for NWS use
> - 1500 UTC today,
> - 1900 UTC today - 0300 UTC tomorrow, 8 hours,
> b) Continuous full-disk test
> - 0315 tomorrow (11:15 pm EDT today) - 1915 UTC (3:15 pm EDT) Saturday, 40 hours
> - 1211-1226 UTC (event1) 1558-1614 UTC (event2) 1858-1914 UTC (event3) today
> c) Outage of GOES-16 ABI products and their delivery for solar calibration [VNIR calibration] and desert monitoring. Note: the times of these events may change due to spacecraft activities.
> 
> NHC RFCs:
> 1. On March 7, ECMWF started disseminating medium-range probabilistic forecasts to its users 40 minutes earlier than before. NHC has requested that the ECMWF ensemble jobs run on WCOSS be moved up correspondingly to enable them to use the latest ensemble run a cycle earlier in their forecast process. The output of these jobs is not disseminated to the public (NHC is the only user outside of NCEP). To be implemented on Tuesday, April 25 at 1930Z (3:30 pm EDT).
> 
> 2. Implement N-AWIPS v7.4.0. This includes: updated NHC cone radii for 2017, NHC updates for 2017 (bug fixes, seasonal updates); fixed missing 2nd tier marine zone; SFMAP performance improvement; fixed PROB2CAT bugs that occurs while composing convective outlook and when lines are drawn outside the international boundaries; bug fix in the ensemble diagnostic function (ENS_CVAL); and other changes. A full list of changes will be provided upon request. To be implemented on Tuesday, April 25 at 1200Z (8 am EDT).
> 
> 




More information about the Ncep.nhc.nco_contacts mailing list